Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4b)
gherardini - 23:00 Thursday 16 May 2024 (64273) Print this report
Operator Report - Afternoon shift
This afternoon science mode stopped: 

14:43UTC for adjusting:
- SQB1 shutter closing (370000 step forward);
- WI CO2 chiller temperature adjustment (#64274);

the ITF unlocked at 14:52UTC during the adjusting time, it relocked at the second attempt just in time to start with the weekly calibrations:
- 16:15UTC: Measurement of checkHrec, NE,WE,BS,PR,SR optical responses, and sensitivity (CALIBRATED_DF_SENSITIVITY);
- 17:14UTC: Measurement of actuators response for NE,WE actuators and NI,WI,BS marionettes (CALIBRATED_DF_PCAL);
- 17:48UTC: Measurement of the TF between CAL and Sc channels (CALIBRATED_DELAY);
- 17:57UTC: Measurement of actuators response for PR and BS mirrors (CALIBRATED_PRWI);
- 19:24UTC: Measurement of actuators response for NI,WI mirrors (CALIBRATED_DF_INtoEND_LN), ITF unlocked opening the NI,WI actuators (DF_LN1calib_to_LN2ready state in CALI node);

the last two relocks worked at the first attempt, science mode started at 20:46UTC.
Images attached to this report
Comments to this report:
bersanetti - 11:49 Monday 20 May 2024 (64301) Print this report

The unlock at 19:24UTC during Calibration is due to the fact that now we actually skip the LowNoise1 actuators configuration and we move directly from HighPower to LowNoise2; therefore acting on the IM relays in LOW_NOISE_1 will kill the actuators in use. Also I am not sure if such calibration measurement is meaningful if we lock with the IM actuators.

The best option in my opinion is to force the switch to LowNoise1 when we do this Calibration measurement so to reinstate the old situation; however this is not straightforward as we anticipate part of the EM configuration during the acquisition to avoid the now unnecessary toggle between coil pairs to avoid segments with half the dynamic.

rolland - 15:27 Tuesday 21 May 2024 (64315) Print this report

After discussion, we agreed that the measurement is still meaningful to transfer the calibration from the PCal to the NI,WI mirror actuator in HP mode (then used to calibrate the BS +PR,SR mirror actuators). Doing the calibration in the new LOW_NOISE_1 mode looks ok (with the warning that calibration analysis should use the CAL_NI,WI_MIR_Z_CORR channels and not the Sc_NI,WI_MIR_Z_CORR channels).

Now, the CALI state DF_LN1calib_to_LN2ready is mainly doing nothing, instead of opening the NI,WI relays as it was done until now, since the NI,WI mirrors are now still used for the ITF control at this step. To be checked during the next calibration data taking that ITF does not unlock anymore after this measurement.

    def main(self):        ezca.put_index(self.index)'Setting the NI,WI actuators to their standard state for LOW_NOISE ITF locks')        #self.message = "WI and NI coil relays set to 10 kOhm series resistor"        self.message = "WI and NI coil relays kept in HP mode"        #sb_10K()        self.timer['SAT_RELAY'] = 1 #ramp of 5 s
Search Help